OpenStudy (anonymous):

PLEASE HELP ME!!!! Compare the function with the parent function. Without graphing, what are the vertex, axis of symmetry, and transformations of the parent function?

OpenStudy (anonymous):

A)(–1/4, –7); x = –1/4; translated to the right 1/4 units and up 7 units B)(1/4, 7); x = 1/4; translated to the right 1/4 units and up 7 units C)(–1/4, 7); x = –1/4; translated to the left 1/4 units and up 7 units D)(1/4, –7); x = 1/4; translated to the right 1/4 units and down 7 units